# Generate-on-Graph: Treat LLM as both Agent and KG in Incomplete Knowledge Graph Question Answering

### v1

@@ 5 Generate-on-Graph (GoG)
In this section, we introduce our method Generate-on-Graph (GoG), which can integrate the knowledge of KGs and LLMs, as well as utilize the reasoning ability of LLMs. The comparison between GoG and other previous methods is illustrated in Figure 3.
LLM as Agent. Motivated by ReAct Yao et al. (2023), we consider the LLM as an agent interacting with an environment to solve tasks. As shown in Figure 3 (d), for each step $i$, the agent first generates a thought $t_{i}\in\mathcal{L}$, where $\mathcal{L}$ is the language space, to decompose the original question (Thought 1) or decide which next sub-question should be solved (Thought 2). Then, based on the thought $t_{i}$, the agent generates an action $a_{i}\in\mathcal{A}$, where $\mathcal{A}$ is the action space, to search information by calling graph database API (Act 1, 2) or generate more information by reasoning and inherent knowledge (Act 4).
Action Space. GoG uses the selecting-generating-answering framework, which consists of three main actions: Search, Generate and Answer.
1. Search[target entity], which aims to find the most relevant top K neighbors of the target entity $e$. It is possible to search multiple entities at one time (Act 3 in Figure 3), for simplicity, we only consider searching one entity at one search action here. As there could be numerous neighbor entities of the target entity. We use the same strategy in ToG Sun et al. (2023) to filter unrelated neighbor entities. First, we retrieve all neighbor relations $r\in{R^{e}}$ linking to the target entity. This process is completed by pre-defined SPARQL queries. Then, we utilize the LLMs to select the top K relations that most related to the current sub-question (last thought). As shown in Act 2 of Figure 3 (d), given the target entity Cupertino, LLMs select the two relation {Located_in, Adjoin} from all neighbor relations. In the end, triples {(Cupertino, located_in, California), (Cupertino, adjoin, Palo Alto)} are appended to context as Obs 3.
2. Generate[sub-question], which make the agent generate new factual triples based on retrieval information and inherent knowledge. As shown in Act 4 of Figure 3 (d), although there is no triple directly representing the timezone of Cupertino, the agent can still infer the timezone of Cupertino based on {(Cupertino, located_in, California), (California, timezone, Pacific Standard Time)}. This process is similar to knowledge graph completion (KGC) Wang et al. (2017). It is also possible for the agent directly generates an entity, which is not explored before, based on its inherent knowledge. Therefore, we have to link the entity to its corresponding MID in the KG. This entity linking process is divided into two steps: (1) We retrieve some similar entities and their corresponding types based BM25 scores. (2) We utilize the LLM to select the most relevant entity based on the types.
3. Finish[answer], indicates that the agent finishes the task with answer. It should be noticed that the agent would also generate "Finish[unknown]", which means that there is not enough information for the agent to answer the question. In this case, we would roll back and search one more hop neighbors of the last target entity.
{tblr}
width = 0.8rows=rowsep=1pt, colspec = Q[452]Q[106]Q[125]Q[106]Q[125], row1 = c, row2 = c, row7 = c, column3 = c, column5 = c, cell12 = c=20.23, cell14 = c=20.23, cell21 = c=50.914, cell32 = c=20.23c, cell34 = c=20.23c, cell42 = c=20.23c, cell44 = c=20.23c, cell52 = c=20.23c, cell54 = c=20.23c, cell62 = c=20.23c, cell64 = c=20.23c, cell71 = c=50.914, cell82 = c, cell84 = c, cell92 = c, cell94 = c, cell102 = c, cell104 = c, cell112 = c,fg=MineShaft, cell113 = fg=MineShaft, cell114 = c,fg=MineShaft, cell115 = fg=MineShaft, cell122 = c, cell124 = c, cell132 = c, cell134 = c, cell142 = c, cell144 = c, cell152 = c, cell154 = c, cell162 = c, cell164 = c, cell172 = c, cell174 = c, vline2 = 1,3-6, vline4 = 1,3-6, vline2,4 = 8,10-11,13-17, hline1-3,7-10,12-13,18 = -, Method & CWQ WebQSP
without external knowledge
@@
RoG w/o planning 43.0 66.9
with external knowledge
CKG IKG-1 CKG IKG-1
w training
RoG 64.5 51.1 88.7 70.1
ChatKBQA 76.5 35.0 78.1 42.6
w/o training
KB-BINDER - - 50.7 34.4
StructGPT - - 76.4 53.7
ToG 47.2 36.7 76.9 60.6
GoG w/GPT-3.5 (Ours) 55.7 43.4 78.7 64.9
GoG w/GPT-4 (Ours) 75.2 61.0 84.4 74.8
@@
Table 1: The Hits@1 scores of different models over two datasets in different settings (%). CKG and IKG-1 denote using complete and incomplete KG (dropping one crucial triple for each question), respectively. We use GPT-3.5 as backbone of KB-BINDER, StructGPT and ToG. Results of the other baselines were re-run by us, more details can be found in Appendix C. The boldface indicates the best result in the same setting.
@@ 6.1 Experiments Setup
The baselines we compare can be divided into two groups: (1) LLM only methods, including standard prompting (IO prompt) Brown et al. (2020), Chain-of-Thought (CoT) prompting Wei et al. (2023) and Self-Consistency (SC) Wang et al. (2023b). (2) Semantic Parsing (SP) methods, including KB-BINDER Li et al. (2023c) and ChatKBQA Luo et al. (2023a). (3) Retrieval Augmented (RA) methods, including StructGPT Jiang et al. (2023), RoG Luo et al. (2023b) and ToG Sun et al. (2023), where RoG is the SOTA among all models requiring fine-tuning. All these methods are evaluated in both complete incomplete KGs.
@@
We use two LLMs as the backbone in our experiments: GPT-3.5 and GPT-4. We do not use Llama-2-70b-chat-hf Touvron et al. (2023), as we found that Llama’s output sometimes did not follow the format we defined in few-shot. We use OpenAI API to call GPT-3.5 and GPT-4. The maximum token length for each generation is set to 256. The temperature parameter is set to 0.7 in all experiments. We use 3 shots in GoG prompts for all the datasets. The prompts we use are listed in Appendix B. The details of re-running other baselines can be found in Appendix C.
{tblr}
width = colspec = Q[220]Q[180]Q[180]Q[180]Q[180], columneven = c, column3 = c, column5 = c, hlines, & IKG-1 IKG-2 IKG-3 IKG-4
CWQ 2.3 4.3 5.9 6.8
WebQSP 2.1 3.6 4.6 5.4
Table 2: The average number of edges deleted under different incompleteness degrees.
@@
For each dataset, we generate four incomplete KGs with varying degrees of completeness: IKG-1/2/3/4, representing randomly dropping 1/2/3/4 crucial triples in the KG. In addition to the crucial triples themselves, all relations between these two entities will also be deleted. The statistics of these incomplete KGs are shown in Table 2. Besides, we also ensure that after deleting these crucial triples, the number of neighbor nodes of the topic entities will not be zero, more details can be found in Appendix D.
      Method       WebQSP
      CKG       IKG-1       IKG-2       IKG-3       IKG-4
      StructGPT       76.4       52.5       49.1       47.9       46.4
      ToG       76.9       60.6       57.8       58.0       57.3
      GoG       78.7       62.2       60.9       59.6       57.4
      CWQ
      CKG       IKG-1       IKG-2       IKG-3       IKG-4
      ToG       47.2       36.7       33.1       32.2       31.7
      GoG       55.7       41.6       37.2       36.7       36.5
Table 3: The Hits@1 scores of prompt based methods (w/ GPT-3.5) under different numbers of missing triples (%). CKG represent using the complete KG. IKG-1/2/3/4 represent randomly dropping 1/2/3/4 crucial triples in the KG.
@@ 6.2 Main Results
Table 1 shows the Hits@1 scores of GoG and all baselines on two datasets in different settings. From the table, we can find that, compared with other prompt based methods, GoG can achieve the state-of-the-art performance on CWQ and WebQSP in both complete and incomplete KG settings.
In the CKG setting, the excellent performance of GoG mainly comes from its dynamic subgraph expansion strategy, which performs better than ToG in question involving compound value types (CVTs), an example of CVT is demonstrated in Figure 4. ToG is likely to think CVT nodes are not worthy to further explore and ignore them, as they do not offer information directly. Our GoG can easily solve this problem by expanding subgraph dynamically, that means if there is not enough information provided by the current subgraph, GoG would search one more hop, so the neighbors of CVT nodes is taken into consideration in this way.
In the IKG setting, the performance of SP methods, such as ChatKBQA and KB-BINDER, significantly declines. This is expected, as these SP methods don’t interact with the KGs, which means they have no idea of the absence of some triples. The performance of RoG, StructGPT and ToG also drop significantly. The performance of ToG and StructGPT on IKG is even worse than that without KG (IO prompt and CoT). That means, these methods still play a role of parsing to find answers rather than effectively integrating internal and external knowledge sources. Besides, in the IKG setting, these methods are even likely to retrieve wrong or unrelated paths which disturb LLMs. Our GoG can alleviate the problem by using the generate action, which utilizes the LLM to generate new factual triples when no direct answer is found after multiple rounds of searches on the KGs. Detailed analysis of answers generated by GoG can be checked in Appendix E.
6.3 Performance under Different Numbers of Missing Triples
In order to explore the impact of different degrees of KG incompleteness on different methods, we evaluate the performance of methods (w/ GPT-3.5) under different numbers of missing triples, the results are demonstrated in Table 3.
It can be found that our GoG outperforms other prompt based methods consistently in different numbers of missing triples. Especially on the CWQ dataset, our GoG has a significant improvement on Hits@1 score, achieving average 7.5% improvement under all settings. That emphasizes the importance of integrate the external and inherent knowledge of LLMs. On the contrary, the performance of ToG on IKG is even much lower than that without KG, which indicates the performance of ToG still depends heavily on the completeness of KGs.
Furthermore, even though the majority of questions in the WebQSP dataset are single-hop questions, GoG can still demonstrate its advantages and perform better than ToG and StructGPT. This is because GoG can leverage the neighboring information of the topic entities to predict the tail entities while other methods can not make full use these information. More details can be found in Appendix F.
@@ 6.4 Performance with Different LLMs
Method WebQSP
CKG IKG-1 NKG
GoG w/GPT-3.5 78.7 64.9 62.2
GoG w/GPT-4 84.4 74.8 65.6
CWQ
CKG IKG-1 NKG
GoG w/GPT-3.5 55.7 43.4 38.8
GoG w/GPT-4 75.2 61.0 55.6
Table 4: The Hits@1 scores of GoG using different backbone models (%). CKG, IKG and NKG denote using complete, incomplete and no KG.
Figure 5: The Hits@1 scores of GoG with different generation strategies on the CWQ (a) and WebQSP (b) (%), w subgraph and w/o subgraph represent generating new factual triples with and without explored subgraph as context in the Generate action, respectively.
Figure 6: The Hits@1 scores of GoG with different number of related triples in the Generate action (%).
@@
In addition, we can also find that, the more powerful the backbone model, the better it utilizes the incomplete KGs. For example, on the WebQSP dataset, the performance of GoG (w/GPT-3.5) using incomplete KG is only 2.5% higher than that without using KG, while the improvement increases to 9.2% when using GPT-4 as backbone.
@@ 6.5 Ablation Study
The Effect of Explored Subgraphs
To explore the influence of explored subgraphs on the performance of GoG, we conduct experiments under different two Generate action (introduced in section 5) strategies: (1) Utilizing explored subgraphs, which is obtained in Search action, as context to generate new factual triples. (2) Generating new factual triples directly without explored subgraphs. As shown in Figure 5, GoG’s performance improves significantly with the help of explored subgraphs. This implies that, even incomplete KGs don’t provide answers directly, the related knowledge they provide is still helpful for models generating corresponding knowledge. There are two potential reasons: (1) Utilizing related subgraphs as context can activate LLMs’ memory corresponding to this knowledge. (2) LLMs can reason new factual triples based on the related subgraphs, as shown in Figure 3 (d).
@@
Aiming to find out how many related triples is required in the Generate action, we perform additional to find out how the number of related triples effect GoG’s performance. We select the most relevant k triples based on BM25. The results are shown in Figure 6. It can be observed that, GoG’s performance first increases and then decreases as the number of related triples increases. This is mainly because noisy and unrelated knowledge are introduced when the number of related triples is large. How to filter valuable knowledge from the explored subgraph could be a future direction.

### v2

@@ 5 Generate-on-Graph (GoG)
In this section, we introduce our method Generate-on-Graph (GoG), which can integrate the knowledge of KGs and LLMs, as well as utilize the reasoning ability of LLMs. The workflow of GoG is illustrated in Figure 3 (c). GoG utilizes the Thinking-Searching-Generating framework, which consists of three main steps: Thinking, Searching and Generating.
5.1 Thinking
Motivated by ReAct Yao et al. (2023), we consider the LLM as an agent interacting with an environment to solve tasks. GoG use the Thinking-Searching-Generating framework to answer questions. As shown in Figure 3 (c), for each step $i$, GoG first generates a thought $t_{i}\in\mathcal{L}$, where $\mathcal{L}$ is the language space, to decompose the original question (Thought 1), decide which next sub-question should be solved (Thought 2) or determine whether it has adequate information to output the final answers (Thought 4). Then, based on the thought $t_{i}$, GoG generates an action $a_{i}\in\mathcal{A}$, where $\mathcal{A}$ is the action space, to search information from the KG (Action 1, 2) or generate more information by reasoning and internal knowledge (Action 3).
Method CWQ WebQSP
w.o. Knowledge Graph
@@ 5.1 Thinking
CKG IKG CKG IKG
w.t. Knowledge Graph / Fine-tuned
RoG Luo et al. (2023) 66.1 54.2 88.6 78.2
ChatKBQA Luo et al. (2024) 76.5 39.3 78.1 49.5
w.t. Knowledge Graph / Not-Training (GPT-3.5)
KB-BINDER Li et al. (2023c) - - 50.7 38.4
StructGPT Jiang et al. (2023) - - 76.4 60.1
ToG Sun et al. (2023) 47.2 37.9 76.9 63.4
GoG (Ours) 55.7 44.3 78.7 66.6
w.t. Knowledge Graph / Not-Training (GPT-4)
ToG Sun et al. (2023) 71.0 56.1 80.3 71.8
GoG (Ours) 75.2 60.4 84.4 80.3
@@
Table 1: The Hits@1 scores of different models over two datasets under different settings (%). CKG and IKG denote using complete and incomplete KG (IKG-40%), respectively. Results of the other baselines were re-run by us  [^2] The evaluation strategy we use differs from that of ToG, which makes the performance of ToG vary from those reported. Further details are available in Appendix D.. The boldface indicates the best result.
5.2 Searching
The search action is invoked by GoG in the form of $Search[e_{i}]$, where $e_{i}$ is the target entity, as illustrated in Action 1 and 2 in Figure 3 (c). While it is possible to search multiple target entities, like $Search[e_{i}^{1},e_{i}^{2},\ldots]$, for simplicity, we only consider searching for one target entity here. This action aims to find the most relevant top-k entities $E_{i}$ from the neighboring entities of the target entity $e_{i}$ based on the last thought $t_{i}$. The search action consists of two steps: Exploring and Filtering.
  • Exploring GoG first uses predefined SPARQL queries to obtain all the relations $R_{i}$ linked to the target entity $e_{i}$. For example, in Figure 3 (c), $e_{1}$={Apple Inc} $R_{1}$={founder, headquarter, CEO}.
  • Filtering After retrieving the relation set $R_{i}$, LLMs are utilized to select the most relevant top-N relations $R^{\prime}_{i}$ based on the last thought $t_{i}$. The prompt used for this step is detailed in Appendix C. In the case of Figure 3 (c), LLMs select $R^{\prime}_{1}$={headquarter} from $R_{1}$={founder, headquarter, CEO} to answer the thought $t_{1}$I need to find out where is the Apple’s headquarters".
Finally, we obtain the most relevant entity set $E_{i}$ based on the target entity $e_{t}$ and the relevant relation set $R^{\prime}_{i}$. As shown in Figure 3 (c), the Observation in step one is {(Apple Inc, headquarter, Cupertino)}, which is attached to the context to enable GoG to generate the next thought.
5.3 Generating
When there is no direct answer from previous Observation, the Generate Action is invoked by GoG in the form of $Generate[t_{i}]$, where $t_{i}$ is the last thought, as illustrated in Action 3 in Figure 3 (c). This action tries to utilize the LLM to generate new factual triples based on retrieval information and internal knowledge. There are three steps in each Generate Action: choosing, generating and verifying.
  • Choosing To provide LLMs some relevant information to generate more accurate triples, we use BM25 Robertson and Zaragoza (2009) to retrieve the most relevant triples from previous Observation. For example, in the Action 3 in Figure 3 (c), we choose {(Cupertino, located_in, California), (Cupertino, adjoin, Palo Alto)} from Observation 1 and 2 as the relevant triples used in LLM generating new triples.
  • Generating After retrieving relevant triples, LLMs are utilized to generate new factual triples based on these relevant triples and their internal knowledge. The generating process will be repeated $n$ times to minimize error and hallucination. As shown in Action 3 of Figure 3 (c), given relevant triples, LLMs generate {(Cupertino, timezone, Pacific Standard Time)} in generated $t_{1}$.
  • Verifying In the end, we use LLMs to verify the generated triples and choose those are more likely to be accurate as the Observation, the prompt used here is shown in Appendix C. As shown in Observation 3 of Figure 3 (c), LLMs only remain {(Cupertino, timezone, Pacific Standard Time)} from all generated triples.
It is also possible for the LLMs to generate an entity that is not explored before. Therefore, we have to link the entity to its corresponding Machine Identifier (MID) in the KG. This entity linking process is divided into two steps: (1) We retrieve some similar entities and their corresponding types based BM25 scores. (2) We utilize the LLM to select the most relevant entity based on the types, the prompt we use is demonstrated in Appendix C.
GoG repeats the above three steps until it obtains adequate information, and then outputs the final answer in the form of $Finish[e_{a}]$, where $e_{a}$ represents the answer entity. It should be noticed that the agent could also generate "$Finish[unknown]"$, which means that there is not enough information for the agent to answer the question. In this case, we would roll back and search one more hop neighbors of the last target entity.
Method CWQ
CKG IKG-20% IKG-40% IKG-60% IKG-80%
ToG 47.2 40.5 37.9 33.7 31.4
GoG 55.7 44.9 44.3 36.2 34.4
WebQSP
CKG IKG-20% IKG-40% IKG-60% IKG-80%
StructGPT 76.0 67.8 60.1 51.7 43.7
ToG 76.9 70.3 61.4 60.6 55.9
GoG 78.7 70.8 66.6 62.6 56.5
Table 2: The Hits@1 scores of prompt based methods (w/ GPT-3.5) under different numbers of missing triples (%). CKG represents using the complete KG. IKG-20%/40%/60%/80% represent randomly drop 20%/40%/60%/80% crucial triples for each question.
Method CWQ
CKG IKG-40% NKG
GoG w/GPT-3.5 55.7 44.3 38.8
GoG w/Qwen-1.5 63.3 49.2 47.0
GoG w/Llama-3 59.6 54.6 54.0
GoG w/GPT-4 75.2 60.4 55.6
WebQSP
CKG IKG-40% NKG
GoG w/GPT-3.5 78.7 66.6 62.6
GoG w/Qwen-1.5 77.9 70.2 65.1
GoG w/Llama-3 77.4 74.4 70.8
GoG w/GPT-4 84.4 80.3 75.7
Table 3: The Hits@1 scores of GoG using different backbone models (%). CKG, IKG-40% and NKG denote using complete, incomplete and no KG. Qwen-1.5 and Llama-3 represent Qwen-1.5-72b-chat and Llama-3-70b-Instruct, respectively.
@@ 6.1 Experiments Setup
The baselines we compare can be divided into three groups: (1) LLM only methods, including standard prompting (IO prompt) Brown et al. (2020), Chain-of-Thought (CoT) prompting Wei et al. (2023) and Self-Consistency (SC) Wang et al. (2023). (2) Semantic Parsing (SP) methods, including KB-BINDER Li et al. (2023c) and ChatKBQA Luo et al. (2024). (3) Retrieval Augmented (RA) methods, including StructGPT Jiang et al. (2023), RoG Luo et al. (2023) and ToG Sun et al. (2023), where RoG is the SOTA among all models requiring fine-tuning.
@@
We use four LLMs as the backbone in our experiments: GPT-3.5, GPT-4, Qwen-1.5-72B-Chat Bai et al. (2023) and LLaMA-3-70B-Instruct Touvron et al. (2023). We use OpenAI API to call GPT-3.5 and GPT-4  [^3] The specific versions of GPT-3.5 and GPT-4 are gpt-3.5-turbo-0613 and gpt-4-0613.. The maximum token length for each generation is set to 256. The temperature parameter is set to 0.7. We use 3 shots in GoG prompts for all the datasets. The prompts we use are listed in Appendix C.
@@
For each dataset, we generate four incomplete KGs with varying degrees of completeness: IKG-20%/40%/60%/80%, representing randomly drop 20%/40%/60%/80% crucial triples for each question. In addition to the crucial triples themselves, all relations between these two entities will also be deleted. The statistics of these IKGs can be found in Appendix E.
@@ 6.2 Main Results
Table 2 shows the Hits@1 scores of GoG and all baselines on two datasets under different settings. From the table, we can find that, compared with other prompt based methods, GoG can achieve the state-of-the-art performance on CWQ and WebQSP under both complete and incomplete KG settings.
Under the CKG setting, the main reasons our GoG outperforms ToG are: (1) GoG decompose the problem into sub-problems each step and focuses on the information needed for each sub-problem during the search process, whereas ToG lacks overall planning, making it prone to repetitive exploration or getting lost during the search. (2) GoG adopts a dynamic subgraph expansion search strategy, while ToG only explores some paths. Therefore, the relevant information obtained in GoG is richer. Moreover, this strategy can better handle compound value types (CVTs), as detailed in Appendix F. A case study is shown in Appendix H.1.
Under the IKG setting, the performance of SP methods significantly declines. This is expected, as these SP methods don’t interact with the KGs, which means they have no idea of the absence of some triples. The performance of ToG and StructGPT on IKG is even worse than that without KG, indicating that these methods still play a role of finding answers rather than effectively integrating internal and external knowledge sources. Our GoG mitigates this issue by using the Generate Action, which utilizes the LLM to generate new factual triples when no direct answer is found. A case study illustrating this is provided in Appendix H.2, and a detailed analysis of the answers generated by GoG can be found in Appendix G.
Figure 4: The Hits@1 scores of GoG with different number of related triples in the Generate Action on the CWQ (a) and WebQSP (b) (%). The backbone LLM is Qwen-1.5-72b-chat.
6.3 Performance under Different Degrees of KG Incompleteness
To investigate how different degrees of KG incompleteness affect different methods, we evaluate the performance of methods (w/ GPT-3.5) under KGs with varying degrees of incompleteness, the results are demonstrated in Table 2.
It can be found that our GoG outperforms other prompt based methods in different degrees of incompleteness. Especially on the CWQ dataset, our GoG has a significant improvement on Hits@1 score, achieving average 5.0% improvement. That emphasizes the importance of integrate the external and internal knowledge of LLMs under incomplete KGs. On the contrary, the performance of ToG on IKG-40% is even lower than that without KG, indicating the performance of ToG still depends heavily on the completeness of KGs.
Even though the majority of questions in the WebQSP dataset are single-hop questions, GoG still outperforms ToG and StructGPT. This is because GoG can leverage the neighboring information of the topic entities to predict the tail entities while other methods can not make full use these information, a case study is shown in Appendix H.2.
@@ 6.4 Performance with Different LLMs
Additionally, we observe that under the NKG setting, Llama-3 consistently outperforms Qwen-1.5, whereas under the CKG setting, the opposite is true. This suggests that the proficiency of LLM as a KG and as an agent is not entirely equivalent. Exploring how different LLMs can leverage their strengths in playing specific roles could be a direction for future research.
Method CWQ
CKG IKG-40%
GoG w.o. Generate 62.7 48.6
GoG w.t. Generate 63.3 50.6
WebQSP
CKG IKG-40%
GoG w.o. Generate 74.7 69.4
GoG w.t. Generate 77.9 71.1
Table 4: The Hits@1 scores of GoG w.t./w.o. Generate Action (%).
@@ 6.5 Ablation Study
We perform additional experiments to find out how the number of related triples effect GoG’s performance. We select the top-k relevant triples based on BM25, as shown in Figure 4. The results indicate that: (1) GoG’s performance significantly improves with relevant subgraphs, likely because these subgraphs activate LLMs’ memory to generate more accurate triples and enable reasoning of new factual triples based on these subgraphs. (2) In most cases, performance initially increases and then decreases as the number of related triples grows. This decline is mainly due to the introduction of noisy and unrelated knowledge.
The Effect of Generate Action
We investigate the effect of the Generate Action, as shown in Figure 4. GoG’s performance is lower without Generate Action, confirming the effectiveness of Generate Action. However, GoG without Generate Action still achieves competitive results because it becomes a pure exploring agent, leading to two outcomes: (1) No false negatives, as all answers come from KGs, and (2) It thoroughly searches KGs for answers, whereas GoG with Generate Action may determine to invoke Generate Action instead of continuing the search.